OpenStudy (anonymous):

use quadratic formula 2x^2-x=-1

OpenStudy (anonymous):

X = -1/2 AND X = 1

OpenStudy (anonymous):

i think that yu are suppose tu add the 1 tu the other side tu make the equation equal 0

OpenStudy (anonymous):

do u need steps??

OpenStudy (anonymous):

for quaadratic equation of form ax^2 + bx + c = 0 roots by quadratic formula are got as follows: -b +/- root(b^2 - 4ac) x = ----------------------- 2a

OpenStudy (anonymous):

this equation has no real roots have u familiar with complex numbers donnice?

OpenStudy (anonymous):

but what wud the solutions be in terms of i

OpenStudy (anonymous):

@jimmy why will it not hv real roots????

OpenStudy (anonymous):

it has x = 1 and x= -1/2

OpenStudy (anonymous):

http://www.mathway.com/problem.aspx?p=algebra this helps a great deal with figuring things out.

OpenStudy (anonymous):

oh - i'll check again hark

OpenStudy (anonymous):

@jimmy i saw my mistake, i missed the minus sign in front of 1 on the RHS. SORRY!!

OpenStudy (anonymous):

it changes to 2x^2 - x + 1 = 0 right?

OpenStudy (anonymous):

oh ok!!

OpenStudy (anonymous):

its roots will be 1 + 7i 1-7i ------ and ------ 4 4 Right jimmy????

OpenStudy (anonymous):

maybe donnice made a typo with the question - can u check it out donnice?

OpenStudy (anonymous):

right hark

OpenStudy (anonymous):

thanks harkirat

OpenStudy (anonymous):

U R WELCOME......